### v1

# Paired Disjunctive Domination Number of Middle Graphs

Abstract
The concept of domination in graphs plays a central role in understanding structural properties and applications in network theory. In this study, we focus on the paired disjunctive domination number in the context of middle graphs, a transformation that captures both adjacency and incidence relations of the original graph. We begin by investigating this parameter for middle graphs of several special graph classes, including path graphs, cycle graphs, wheel graphs, complete graphs, complete bipartite graphs, star graphs, friendship graphs, and double star graphs. We then present general results by establishing lower and upper bounds for the paired disjunctive domination number in middle graphs of arbitrary graphs, with particular emphasis on trees. Additionally, we determine the exact value of the parameter for middle graphs obtained through the join operation. These findings contribute to the broader understanding of domination-type parameters in transformed graph structures and offer new insights into their combinatorial behavior.
keywords:
Paired disjunctive domination, middle graphs, trees, graph operations.
volume: 185issue: 1
Paired Disjunctive Domination Number of Middle Graphs
1 Introduction
Graph theory, a cornerstone of discrete mathematics, offers a robust framework for modeling and analyzing intricate systems. Its wide-ranging applications include domains such as computer science, biology, transportation, and the social sciences, where graphs provide an intuitive means of representing real-world structures—vertices symbolize individual elements, while edges denote interactions or connections among them. Due to this representational power, graph theory has become a crucial tool in addressing complex challenges, from optimizing communication infrastructures to understanding chemical compound structures.
Within this vast field, domination theory has emerged as a particularly rich area of study. It explores how specific subsets of vertices can exert control or influence over the entire graph. Domination-related parameters play a critical role in both theoretical investigations and practical scenarios, including network defense, facility deployment, and efficient allocation of limited resources. A noteworthy advancement in this area is the paired disjunctive domination parameter—a more recent refinement that introduces an element of redundancy and robustness into domination strategies. This parameter determines the smallest number of vertex pairs needed such that every vertex in the graph either belongs to one of the selected pairs or is adjacent to at least one vertex from a pair. Such a framework is highly relevant for contexts requiring reliability and fault tolerance, including resilient communication networks and strategic emergency response planning.
A practical example of paired disjunctive domination can be observed in urban emergency response systems. Consider a city where ambulance stations must be strategically positioned to ensure prompt medical assistance in case of emergencies. By placing ambulances at designated locations and ensuring that each region is directly served or adjacent to a station, city planners can achieve efficient coverage while maintaining backup support in case of service disruptions. This type of redundancy is crucial in optimizing response times and enhancing overall service reliability.
A fundamental challenge in modern network design is optimizing the balance between resource allocation and redundancy. Since critical resources are often costly, deploying them uniformly across an entire network is impractical. Moreover, the potential failure of resources at specific nodes necessitates redundancy and backup mechanisms, which, while essential for network reliability, impose additional resource demands. This problem has been extensively studied through graph-theoretic models, wherein researchers identify strategically positioned vertex subsets that maintain network connectivity while satisfying redundancy constraints.
Domination theory, including its extensions such as total domination, disjunctive total domination, paired domination, and more recently, paired disjunctive domination, forms a well-established area in graph theory with numerous applications and a rich body of literature [4, 7, 9, 10, 8]. Foundational surveys by Haynes et al. [4, 7] provide comprehensive insights into these parameters and their theoretical significance.
Given a graph $G=(V,E)$, a subset $S\subseteq V$ is defined as a dominating set if every vertex in $V-S$ has at least one neighbor in $S$. The domination number, denoted $\gamma(G)$, represents the minimum cardinality of a dominating set in G. A subset $S\subseteq V$ is a total dominating set (TDS) if every vertex in $V$ is adjacent to at least one vertex in $S$, and the total domination number, $\gamma_{t}(G)$, is the minimum cardinality of a TDS.
A matching in a graph is a collection of pairwise non-adjacent edges, and a perfect matching is one that includes every vertex exactly once. A paired dominating set (PDS) is a dominating set $S$ for which the subgraph induced by $S$, denoted $G[S]$, has a perfect matching. This ensures that every selected vertex has a designated partner within the set. It is known that any graph without isolated vertices admits a PDS, since the endpoints of a maximal matching satisfy this condition. The paired domination number, $\gamma_{pr}(G)$, represents the minimum size of such a set. This concept was originally introduced by Haynes and Slater [5, 6] in the context of modeling redundancy in security systems, where guards are deployed in mutually supportive pairs.
The concept of the disjunctive domination was introduced by Goddard et al. in 2014 as a variation of original domination concept [3]. In a graph G, a subset $D\subseteq V$ is called a b-disjunctive dominating set (bDD-set) if every vertex not in $D$ either has at least one neighbor in $D$ or has at least $b$ vertices in $D$ at distance two. When $b=2$, the set is referred to as a disjunctive dominating set (2DD-set), and the minimum such set defines the disjunctive domination number $\gamma_{d}(G)$. Various studies have explored algorithmic aspects and bounds of disjunctive domination, including its total version, where the induced subgraph $G[D]$ contains no isolated vertices. A subset $D\subseteq V$ in an isolate-free graph $G=(V,E)$ is called a paired disjunctive dominating set (PDD-set) if it satisfies the following conditions:
  • D is a disjunctive dominating set, meaning every vertex in $V-D$ has at least one neighbor in D or at least two vertices in D at distance two.
  • The subgraph $G[D]$ contains a perfect matching, ensuring that every vertex in D is paired with another vertex via an edge.
The paired disjunctive domination number, denoted as $\gamma_{pd}(G)$, represents the minimum cardinality of a $PDD-set$ in G. A $PDD-set$ of cardinality $\gamma_{pd}(G)$ is called a $\gamma_{pd}$-set [8].
We consider the middle graph, a fundamental structural transformation of a given graph that plays a key role in various theoretical and applied settings in this study. The middle graph $M(G)$ of a graph G is obtained by introducing a new vertex for each edge of G and establishing edges between these new vertices whenever their corresponding edges in G share a common vertex. Middle graphs play a critical role in network topology and interconnection structures, particularly in wireless communication networks, where they help optimize connectivity and signal distribution.
In this paper, we investigate the paired disjunctive domination number of middle graphs. Specifically, we establish lower and upper bounds for this parameter in terms of the order of the graph G. Furthermore, we determined the parameter’s value for the middle graph of graphs obtained under certain graph operations. In addition, we explicitly determine the paired disjunctive domination number for various graph families, including special graphs, double star graphs, and friendship graphs.
The vertex set of middle graph $M(G)$ is defined as $V(M(G))=V(G)\cup\mathcal{M},$ where $V(G)=\{i:1\leq i\leq n\}$ and =$\mathcal{M}=\{e_{ij}:$ $ij\in E(G)\}$ denotes the set of new vertices corresponding to the edges of $G.$ For any two vertices $i,j\in V(G)$, if they are adjacent then the edge $ij\in E(G)$ is represented by a new vertex $e_{ij}$ in the middle graph $M(G).$ Therefore, edge set can be described as $E(M(G))=\{ie_{ij},e_{ij}j:ij\in E(G)\}\cup E(L(G))$, where $L(G)$ is the line graph of $G.$
2 Known Results
In this section, we provide some fundamental results related to the paired disjunctive domination parameter, which was introduced by Henning [8] and further developed by Golpek and Aytac [12, 11]. This will facilitate the flow of the subsequent parts of the paper.
Observation 1
[8] For an isolate-free graph $G,$ $\gamma^{d}(G)\leq\gamma_{pr}^{d}(G)\leq\gamma_{pr}(G)$, and $\gamma^{d}(G)\leq\gamma_{pr}^{d}(G)\leq 2\gamma^{d}(G).$
Observation 2
[12] For an isolate-free graph $G,$ $\gamma_{t}^{d}(G)\leq\gamma_{pr}^{d}(G).$
Observation 3
[11] If $T$ is a tree and $D$ is any $PDD$-set, then every support vertex of $T$ is included in $D$ with a its neighbour or $D$ may contains at least two neigbours of the support vertices with their neighbours.
Theorem 2.1
[8] For paired disjunctive domination number of some special graphs:
  • i)
    $\gamma_{pr}^{d}(C_{n})=2\left\lceil\frac{n}{5}\right\rceil$ with $n\geq 3,$
  • ii)
    $\gamma_{pr}^{d}(P_{n})=2\left\lceil\frac{n+1}{5}\right\rceil$ with $n\geq 2,$
  • iii)
    $\gamma_{pr}^{d}(K_{n})=\gamma_{pr}^{d}(K_{m,n})=2$ with $n\geq 2$.
Theorem 2.2
[11]Let $G$ be isolated-free graph, then $2\leq\gamma_{pr}^{d}(G)\leq n.$
3 Results about Special Graph Structures
In this section, we investigate the paired disjunctive domination parameter in middle graphs of several special graph classes.
Theorem 3.1
Let $G$ be a graph on $n$ vertices and without isolated vertices. Suppose the maximum degree in $G$ satisfies $\Delta(G)=n-1$, then
$\gamma_{pr}^{d}(M(G))=2.$
Proof 3.2
Let the vertex set of $G$ be $V(G)=\{v_{1},v_{2},...,v_{n}\}$, and assume that the vertex with degree $n-1$ is $v_{1}$. Consider the middle graph $M(G)$ and select two subdivision vertices, $v_{12},v_{13}$. Since these two vertices are adjacent, they satisfy the paired property. Moreover, all vertices $M(G)$, except for those adjacent to $v_{12}$ and $v_{13}$, are at most two distances away from these two vertices. Thus, the set $\{v_{12},v_{13}\}$ forms a paired disjunctive domination set (PDD-set) in $M(G)$, which implies that $\gamma_{pr}^{d}(M(G))\leq 2$. Furthermore, by Theorem 2.2, it is known that the paired disjunctive domination number is at least 2. That is, $\gamma_{pr}^{d}(M(G))\geq 2$. Consequently, we conclude that $\gamma_{pr}^{d}(M(G))=2.$
Corollary 3.3
Let $G\simeq H$ where $H\in\{K_{n},K_{1,n},W_{1,n}\}$. Then, as a consequence of theorem 3.1, we have
$\gamma_{pr}^{d}(M(H))=2.$
Theorem 3.4
Let $K_{m,n}$ be complete bipartite graph with $m+n$ vertices. Then
$\gamma_{pr}^{d}(M(K_{m,n}))=2.$
Proof 3.5
It is known that the paired disjunctive domination number is at least 2. Therefore, $\gamma_{pr}^{d}(M(K_{m,n}))\geq 2$. Let us denote the vertices of the complete bipartite graph $K_{m,n}$ as $V_{1}=\{v_{1},v_{2},\ldots,v_{m},v_{m+1},\ldots,v_{m+n}\}.$ In constructing the middle graph, consider any two subdivision vertices adjacent to vertex $v_{1}$, and define the set $S=\{v_{1i},v_{1j}\},$ where $i,j\in\{m+1,m+2,\ldots,m+n\}$. In this case, since all other vertices are within distance 2 of these two vertices, and $v_{1i}$ and $v_{1j}$ are also adjacent to each other, the set $S$ forms a paired disjunctive dominating set for the graph $M(K_{m,n})$. Thus, we obtain $\gamma_{pr}^{d}(M(K_{m,n}))\leq 2,$ which completes the proof.
Theorem 3.6
Let $C_{n}$ be the cycle. Then
$\gamma_{pr}^{d}(M(C_{n}))=2\lceil\frac{n}{4}\rceil.$
Proof 3.7
Let us denote the vertex set of the middle graph $M(C_{n})$ as the union of two disjoint subsets
$V_{1}=\{v_{i}\mid 1\leq i\leq n,\,v_{i}\in V(C_{n})\},\quad V_{2}=\{u_{j}\mid 1\leq j\leq n,\,u_{j}\in SD\},$
that is, $V(M(C_{n}))=V_{1}\cup V_{2}$. We first construct an upper bound for $\gamma_{pr}^{d}(M(C_{n}))$. Define the set
$D=\begin{cases}\{u_{j}\mid 1\leq j\leq n,\ j\equiv 1\text{ or}2\pmod{4}\},n\not\equiv 1(mod4)\\ \{u_{j}\mid 1\leq j\leq n,\ j\equiv 1\text{ or}2\pmod{4}\}\cup\{u_{n},v_{n}\},n\equiv 1(mod4).\end{cases}$
and then $S=D$.
Since every vertex in $M(C_{n})$ is either in $S$ or disjunctively dominated by a vertex in $S$, and since the subgraph induced by $S$ contains a perfect matching, it follows that $S$ is a paired disjunctive dominating set (PDD-set) of $M(C_{n})$ for all cases of $n$ modulo 4. Furthermore, we have $|S|=|D|=2\left\lceil\frac{n}{4}\right\rceil$. Hence,
$\gamma_{pr}^{d}(M(C_{n}))\leq|S|=2\left\lceil\frac{n}{4}\right\rceil.$ (1)
Now, we establish a lower bound for $\gamma_{pr}^{d}(M(C_{n}))$. Let $D^{\prime}$ be a $\gamma_{pr}^{d}$-set of $M(C_{n})$. Consider the matching in subgraph $M(C_{n})[D^{\prime}]$ given by
$M=\{u_{i_{1}}u_{i_{1}+1},u_{i_{2}}u_{i_{2}+1},\dots,u_{i_{r}}u_{i_{r}+1}\},$
where $1\leq i_{1}<i_{1}+1<i_{2}<i_{2}+1<\dots<i_{r}<i_{r}+1\leq n$.
Notice that between any two consecutive matched pairs $u_{i_{k}}u_{i_{k}+1}$ and $u_{i_{k+1}}u_{i_{k+1}+1}$, there are at most two vertices of $M(C_{n})$ that are not in $D^{\prime}$ for any $k\in\{1,\dots,r\}$. Therefore, we obtain
$n\leq 2r+2r=4r,$
which implies that
$|M|=2r\geq 2\left\lceil\frac{n}{4}\right\rceil.$ (2)
Combining the upper bound (1) and lower bound (2), we conclude that $\gamma_{pr}^{d}(M(C_{n}))=2\left\lceil\frac{n}{4}\right\rceil.$
Theorem 3.8
Let $P_{n}$ be the path with $n$ vertices. Then
$\gamma_{pr}^{d}(M(P_{n}))=\begin{cases}2\lceil\frac{n-1}{4}\rceil,&n\equiv 2,3\pmod{4}\\ 2\lceil\frac{n-1}{4}\rceil+2,&\text{otherwise.}\end{cases}$
Proof 3.9
Let us denote the vertex set of the middle graph $M(P_{n})$ as the union of two disjoint subsets:
$V_{1}=\{v_{i}\mid 1\leq i\leq n,\,v_{i}\in V(P_{n})\},\quad V_{2}=\{u_{j}\mid 1\leq j\leq n-1,\,u_{j}\in SD\}.$
, that is, $V(M(P_{n}))=V_{1}\cup V_{2}$. We first construct an upper bound for $\gamma_{pr}^{d}(M(P_{n}))$. Define the set
$D=\{u_{j}\mid 1\leq j\leq n-1,\ j\equiv 1,2\pmod{4}\}.$
Now, consider the following cases:
  • If $n\equiv 0,1\pmod{4}$, let $S=D\cup\{u_{n-1},v_{n}\}$.
  • If $n\equiv 2\pmod{4}$, let $S=D\cup\{v_{n}\}$.
  • If $n\equiv 3\pmod{4}$, then $S=D$.
In all cases, $S$ forms a paired disjunctive dominating set (PDD-set) in $M(P_{n})$. Furthermore, we observe that If $n\equiv 2,3\pmod{4}$, then $|S|=|D|=2\left\lceil\frac{n-1}{4}\right\rceil$. Otherwise, $|S|=|D|+2=2\left\lceil\frac{n-1}{4}\right\rceil+2$. Thus, we obtain the following upper bound:
$\gamma_{pr}^{d}(M(P_{n}))\leq|S|=\begin{cases}2\left\lceil\frac{n-1}{4}\right\rceil,&\text{if }n\equiv 2,3\pmod{4},\\ 2\left\lceil\frac{n-1}{4}\right\rceil+2,&\text{otherwise}.\end{cases}$ (3)
Now we construct a lower bound for $\gamma_{pr}^{d}(M(P_{n}))$. Let $D^{\prime}$ be a minimum PDD-set in $M(P_{n})$. Consider the induced subgraph $M(P_{n})[D^{\prime}]$ and let $M=\{u_{i_{1}}u_{i_{1}+1},u_{i_{2}}u_{i_{2}+1},\dots,u_{i_{r}}u_{i_{r}+1}\}$ be the matching in $M(P_{n})[D^{\prime}]$, where
$11\leq i_{1}<i_{1}+1<i_{2}<i_{2}+1<\dots<i_{r}<i_{r}+1\leq n-$ 1 .
Two cases arise:
  • Case 1: If $n\equiv 2,3\pmod{4}$.
    Note that between any two consecutive matched pairs $u_{i_{k}}u_{i_{k}+1}$ and $u_{i_{k+1}}u_{i_{k+1}+1}$, there are at most two vertices in $M(P_{n})$ that are not in $D^{\prime}$, for each $k\in\{1,\dots,r\}$. Hence, we obtain $n-1\leq 2r+2r=4r$ and therefore
    $|M|=2r\geq 2\left\lceil\frac{n-1}{4}\right\rceil.$ (4)
  • Case 2: If $n\equiv 0,1\pmod{4}$.
    In this case, let $N=M\cup\{v_{n},u_{n-1}\}$. By Case 1, we know that $|M|=2r$, and therefore
    $|N|=2r+2\geq 2\left\lceil\frac{n-1}{4}\right\rceil+2.$ (5)
Combining the lower bounds obtained in both cases (4) and (5) with the previously established upper bounds (3), we conclude that $\gamma_{pr}^{d}(M(P_{n}))=\begin{cases}2\left\lceil\frac{n-1}{4}\right\rceil,&\text{if }n\equiv 2,3\pmod{4},\\ 2\left\lceil\frac{n-1}{4}\right\rceil+2,&\text{otherwise}.\end{cases}$
Theorem 3.10
Let $F_{n}$ be the friendship graph with $n\geq 2$. Then
$\gamma_{pr}^{d}(M(F_{n}))=2.$
Proof 3.11
Let the vertices and edges of friendship graph be labeled by $V(F_{n})=\{v_{i}\mid 1\leq i\leq 2n+1\}$ and $E(F_{n})=\{v_{1}v_{i}|i\in\{2,...2n+1\}\}\cup\{v_{i}v_{i+1}|i=2k\text{ and }k=\{1,2,...,n\}\}$. Let $M(F_{n})$ be the middle graph of the friendship graph $F_{n}$. Let the vertex set of $M(F_{n})$ be $V(M(F_{n}))=V_{1}\cup V_{2}$, where
$V_{1}=V(F_{n}),\quad V_{2}=\{u_{j}\mid 1\leq j\leq 3n,\ u_{j}\in\text{SD}\}.$
Let the central vertex, which is common to all $n$ triangles in $F_{n}$, be denoted by $v_{1}$.
From Theorem 2.2, we know that for any graph $G$, it holds that $\gamma_{pr}^{d}(G)\geq 2$. Therefore,
$\gamma_{pr}^{d}(M(F_{n}))\geq 2.$ (6)
Now, consider the set $S=\{v_{1},u_{i}\}$, where $d_{M(F_{n})}(v_{1},u_{i})=1$. By construction, this set $S$ forms a paired disjunctive dominating set (PDD-set) in $M(F_{n})$, since every vertex is either in $S$ or disjunctively dominated by a vertex in $S$, and the induced subgraph $M(F_{n})[S]$ contains a perfect matching. Hence, we obtain
$\gamma_{pr}^{d}(M(F_{n}))\leq 2.$ (7)
From (6) and (7), proof is completed.
Theorem 3.12
Let $D_{n,m}$ be the double star graph. Then
$\gamma_{pr}^{d}(M(D_{n,m}))=4.$
Proof 3.13
Let the vertices and edges of double star graph be labeled by $V(D(n,m))=\{v_{s}|1\leq s\leq n\}\cup\{v_{t}|n+1\leq t\leq m\}\cup\{v_{n+m+1},v_{n+m+2}\}$ and $E(D_{n,m})=\{v_{n+m+1}v_{n+m+2},v_{s}v_{n+m+1},v_{t}v_{n+m+2}\}$. Let $M(D_{n,m})$ denote the middle graph of the double star graph $D_{n,m}$. Let the vertex set of $M(D_{n,m})$ be partitioned as $V(M(D_{n,m}))=V_{1}\cup V_{2}$, where
$V_{1}=\{v_{i}\mid 1\leq i\leq n+m+2,\ v_{i}\in V(D_{n,m})\},V_{2}=\{u_{j}\mid 1\leq j\leq n+m+1,\ u_{j}\in\text{SD}\}.$
We know from Theorem 2.2 that $\gamma_{pr}^{d}(G)\geq 2$ for any graph $G$. Thus, $\gamma_{pr}^{d}(M(D_{n,m}))\geq 2$. To complete the proof, we now show that no set of two vertices can form a PDD-set in $M(D_{n,m})$, and therefore $\gamma_{pr}^{d}(M(D_{n,m}))\geq 4$.
Assume, for contradiction, that there exists a paired disjunctive dominating set $S^{\prime}$ of $M(D_{n,m})$ with $|S^{\prime}|=2$. Since $S^{\prime}$ must induce a perfect matching and ensure that every vertex in $M(D_{n,m})$ is either in $S^{\prime}$ or disjunctively dominated by a vertex in $S^{\prime}$, the only possibility is that $S^{\prime}$ consists of two adjacent vertices. That is, $S^{\prime}=\{x,y\}$ such that $d_{M(D_{n,m})}(x,y)=1$.
However, due to the structure of the double star graph, the vertices $x$ and $y$ together can disjunctively dominate at most a localized region of the graph either within a single star component or among a small subset of adjacent vertices. This is insufficient to cover the entire graph, especially when $n,m\geq 1$, since the graph contains at least $n+m+3$ vertices in total. Thus, not all vertices are disjunctively dominated by $S^{\prime}$, and $S^{\prime}$ cannot be a PDD-set. Hence, no set of two vertices can form a PDD-set in $M(D_{n,m})$, implying
$\gamma_{pr}^{d}(M(D_{n,m}))\geq 4.$
In the case of $M(D_{n,m})$, we construct a specific set of four vertices that forms a PDD-set and achieves this bound. Let the two central vertices of the stars in $D_{n,m}$ be denoted by $v_{a}$ and $v_{b}$, and let $u_{a}$ and $u_{b}$ be two edge-subdivision vertices adjacent to $v_{a}$ and $v_{b}$ respectively in $M(D_{n,m})$. Consider the set $S=\{v_{a},u_{a},v_{b},u_{b}\}.$ By construction, every vertex in $M(D_{n,m})$ is either in $S$ or disjunctively dominated by a vertex in $S$, and the subgraph induced by $S$ contains a perfect matching. Therefore, $S$ is a paired disjunctive dominating set of $M(D_{n,m})$, and we have
$\gamma_{pr}^{d}(M(D_{n,m}))\leq 4.$
Therefore, we conclude that $\gamma_{pr}^{d}(M(D_{n,m}))=4.$
4 Main Results
In this section, we present general results on the paired disjunctive domination of middle graphs. In particular, we establish lower and upper bounds for trees and general graphs, and provide exact value related to the join operation.
Lemma 4.1
Let $G$ be a connected graph with $n\geq 5$ vertices and $S$ be a paired disjunctive dominating set of $M(G).$ Then there exist $S^{\prime}\subset\mathcal{M}$ a paired disjunctive dominating set of $M(G)$ with $\left|S^{\prime}\right|\leq\left|S\right|.$
Proof 4.2
If $S\subset\mathcal{M}$, then the result is satisfied by choosing $S^{\prime}=S$. Otherwise, the set $S$ also contains vertices from $V(G)$. In this case, let us select a vertex $v\in S\cap V(G)$. If all vertices $e_{iv}\in\mathcal{M}$ (or $e_{vi}$) adjacent to $v$ are in $S$, then two cases arise:
Case 1: If the vertex $v$ does not contribute to the paired property of the set $S$, then we take $S_{1}=S-\{v\}$.
Case 2: If the vertex $v$ contributes to the paired property of the set $S$, then there are two subcases:
  • If all vertices $e_{iv}\in\mathcal{M}$ adjacent to $v$ are in $S$, then we select a vertex $e_{mn}\in\mathcal{M}\setminus S$ from the set $N_{2}(e_{iv},M(G))$, and take $S_{1}=(S\cup\{e_{mn}\})\setminus\{v\}$.
  • If not all adjacent $e_{iv}\in\mathcal{M}$ are in $S$, suppose $e_{iv}\in M\setminus S$. In this case, we take $S_{1}=(S\cup\{e_{iv}\})\setminus\{v\}$.
This process terminates in a finite number of steps, and eventually, a set $S^{{}^{\prime}}$ consisting only of vertices in $\mathcal{M}$ is obtained. Furthermore, in each step, $S^{{}^{\prime}}$ remains a paired disjunctive dominating set. Therefore, $S^{{}^{\prime}}$ is a paired disjunctive dominating set of $M(G)$. Therefore, we construct $|S^{{}^{\prime}}|\leq|S|$.
Lemma 4.3
Let $G$ be a connected graph with $n\geq 5$ vertices. Let $v\in V(G)$ be a vertex that is distinct from a support vertex. Then,
$\gamma_{pr}^{d}(M(G-v))\leq\gamma_{pr}^{d}(M(G))\leq\gamma_{pr}^{d}(M(G-v))+2.$
Proof 4.4
(i) First, let us show that $\gamma_{pr}^{d}(M(G))\leq\gamma_{pr}^{d}(M(G-v))+2$. Let $S$ be a $PDD-set$ of the graph $M(G-v)$. In this case, for every $w\in N_{\leq 2}(v,G)$, we must have either $w\in S$ or $e_{iw}\in\mathcal{M}$, which implies that $e_{iw}\in S$. As a result, for some $w_{1},w_{2}\in N_{G}(v)$, the set $S\cup\{e_{w_{1}v},e_{vw_{2}}\}$ is a $PDD-set$ for $M(G)$. Thus, we obtain the conclusion
$\gamma_{pr}^{d}(M(G))\leq\gamma_{pr}^{d}(M(G-v))+2.$
(ii) Now, let us show that $\gamma_{pr}^{d}(M(G-v))\leq\gamma_{pr}^{d}(M(G))$. Let $S$ be a $\gamma_{pr}^{d}-set$ of $M(G)$. From Lemma 4.1, we assume that $S\subseteq\mathcal{M}$. Let us consider the set
$S_{v}=N_{\leq 2}(v,M(G))\cap S.)$
Since $S$ is a $\gamma_{pr}^{d}-set$, it is easy to see that $|S_{v}|\geq 1$. Let the neighbors and the vertices at distance $2$ from vertex $v$ in the graph $G$ be denoted by $w\in N_{G}(v)$ and $t\in N_{2}(v,G)$, respectively. Assume that $|S_{v}|=k$, where $k\geq 1$, and the vertices in $S_{v}$ are of the form $e_{wv}$ or $e_{tw}$. Based on these types of vertices, two cases arise:
Case 1: If $e_{wv}\in S$, then the set
$S_{1}=(S-e_{wv})\cup\{w\}$
is a $PDD-set$ for the graph $M(G-v)$.
Case 2: If $e_{tw}\in S_{v}$, then the set $S_{1}=(S-e_{tw})\cup\{t\}$ is a $PDD-set$ for the graph $M(G-v)$. For each vertex in $S_{v}$, the corresponding operation described above is applied depending on which of the two cases it falls under, resulting in a $PDD-set$ $S_{k}$ for the graph $M(G-v)$. It is easy to verify that $|S_{k}|=|S|$. Hence, it follows that $\gamma_{pr}^{d}(M(G-v))\leq\gamma_{pr}^{d}(M(G)).$
Theorem 4.5
Let $G$ be any graph of order $n.$ Then
$2\leq\gamma_{pr}^{d}(M(G))\leq\gamma_{pr}^{d}(M(P_{n}).$
Proof 4.6
Let $G\subseteq K_{n}$. By adding the necessary edges to the graph $G$, we can obtain a graph isomorphic to $K_{n}$. Therefore, $M(G)$ becomes a subgraph of $M(K_{n})$. Since any paired disjunctive domination set $(PDD-set)$ of $M(G)$ is also a $PDD-set$ of $M(K_{n})$, it follows that
$\gamma_{pr}^{d}(M(G))\geq\gamma_{pr}^{d}(M(K_{n}))=2.$
This provides the left-hand side of the inequality stated in Theorem 2.1(iii).
Let $G$ be a spanning tree, and let $S$ be a $\gamma_{pr}^{d}$-set of $M(G)$. By Lemma 4.1, we may assume that $S\subseteq\mathcal{M}$. The most vulnerable structure is the path graph considering the number of connections. This implies
$|S|\leq\gamma_{pr}^{d}(M(P_{n})).$
Since $S$ is a $\gamma_{pr}^{d}$-set of $M(G)$, it follows that
$\gamma_{pr}^{d}(M(G))\leq\gamma_{pr}^{d}(M(P_{n})).$
Thus, the desired inequality is obtained.
Theorem 4.7
Let $T$ be a tree graph, which is not a star. Let $SSV=\{v_{i}:1\leq i\leq k\}$ be the strong support vertices of the tree $T.$ For any $u,w\in leaf(T)$, under the conditions that $N_{G}(u)\cap N_{G}(w)=\emptyset$ and $d_{G}(u,w)>3,$ it follows that
$\gamma_{pd}^{d}(M(T))\geq 2k+2(\left|leaf(T)\right|-\sum\limits_{i=1}^{k}(\deg(v_{i})-1)).$
Proof 4.8
Let the set of leaves of the tree $T$ be denoted by $\text{leaf}(T)$. The set $\text{leaf}(T)$ contains strong support vertices. Let us denote these vertices by the set
$\text{SSV}=\{v_{1},\dots,v_{k}\}.$
In this case, the number of support vertices in the graph is given by
$|\text{leaf}(T)|-\sum_{i=1}^{k}(\deg(v_{i})-1).$
The vertices labeled as $e_{iv_{j}}$, where $i=1,\dots,\deg_{T}(v_{j})$ in $M(T)$ corresponding to the edges incident to each vertex $v_{j}$, for $j=1,\dots,k$, form a complete graph with $deg_{G}(v_{j})$ vertices. Let $S$ be a PDD-set of $M(T)$. Then, by Theorem 2.1(iii), it is easy to observe that $|S|\geq 2k$.
If the graph contains a support vertex, then by Observation 3, the set $S$ must include at least one support vertex. Moreover, to maintain the paired property, the neighbors of these support vertices must also be included in $S$. Hence, the following inequality is obtained $\gamma_{pd}^{d}(M(T))\geq 2k+2(\left|leaf(T)\right|-\sum\limits_{i=1}^{k}(\deg(v_{i})-1)).$
Corollary 4.9
Let $T$ be a tree, distinct from a star graph, and containing no support vertices. Then
$\gamma_{pd}^{d}(M(T))\geq 2k.$
Corollary 4.10
Let $T$ be a tree without support vertices, then
$\gamma_{pd}^{d}(M(T))\geq 2(\left|leaf(T)\right|.$
Remark 4.11
If $diam(T)=4$, then equality holds for the corollaries above.
Theorem 4.12
Let $G$ and $H$ be two graphs and $G+H$ be join of them. Then
$\gamma_{pd}^{d}(M(G+H))=2.$
Proof 4.13
Let vertex set of $M(G+H)$ is $V(M(G+H))=V(G)+V(H)\cup\mathcal{M}$ where $\mathcal{M}$ is subdivision vertices of $G+H$, where $V(G)=\{v_{i}:1\leq i\leq n_{1}\}$, $V(H)=\{u_{j}:1\leq j\leq n_{2}\}$, and =$\mathcal{M}=\{e_{v_{i}u_{j}},e_{v_{i}v_{t}},e_{u_{{}_{j}}u_{p}}:i,t=1,...,n_{1},i\neq t,j,p=1,...,n_{2},j\neq p\}$.
Let $S$ be a $PDD-set$ of $M(G+H)$. From Theorem 2.2, it is known that $|S|\geq 2$. Suppose $S=\{e_{v_{i}u_{j}},e_{v_{i}u_{p}}\}$ (or similarly $S=\{e_{v_{i}u_{j}},e_{v_{t}u_{j}}\}$. In this case, all vertices of $V(M(G+H))$ would be disjunctively dominated. Since the vertices in $S$ are adjacent, the set also satisfies the paired property. Thus, $|S|\leq 2$. From the lower and upper bounds, it follows that $|S|=2$.